Find equivalent functional definition for each of the following and draw a graph of a function.


f(x)=2x -|x| + |x-1|

1
Expert's answer
2022-06-10T10:40:02-0400
"x<0, f(x)=2x-(-x)-(x-1)=2x+1"

"0\le x<1, f(x)=2x-x-(x-1)=1"

"x\ge 1, f(x)=2x-x+(x-1)=2x-1"

"f(x)=\begin{cases}
 2x+1, &x<0 \\
 1, & 0\le x<1 \\
2x-1, & x\ge 1
\end{cases}"